OpenStudy (anonymous):

evaluate log2 4-log2 2

OpenStudy (anonymous):

logn a - logn b = logn (a/b) Use that relationship as a first step

OpenStudy (anonymous):

\[\log _{n}a - \log _{n}b = \log _{n}\frac{ a }{ b }\]

OpenStudy (anonymous):

What do you get when you apply that?

OpenStudy (anonymous):

i don't even know how to do any of this

OpenStudy (anonymous):

You don't see how your problem matches up to the equation I wrote? What will your log2 4-log2 2 equal...looking at that next to my equation?

OpenStudy (anonymous):

 What goes in the box?